# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ 4.1 Quasiconformal Mappings
Quasiconformal mappings generalize conformal mappings by extending orientation-preserving homeomorphisms to those with bounded conformality distortions. Mathematically, suppose $\Omega$ is a domain in $\mathbb{C}$, a mapping $f\colon\Omega\rightarrow\mathbb{C}$ is a quasi-conformal map if it satisfies the Beltrami equation
@@
for some measurable complex-valued function $\mu$ with $\|\mu\|_{\infty}<1$, where $\frac{\partial f}{\partial z}=\frac{1}{2}(f_{x}-if_{y})$ and $\frac{\partial f}{\partial\bar{z}}=\frac{1}{2}(f_{x}+if_{y})$. $\mu$ is called the complex dilation or Beltrami coefficient (BC) of $f$, which measures the local deviation from a conformal map.
Let $f\colon(x,y)\rightarrow(u,v)$ and $\mu=\rho+i\tau$, and denote by $\mathbf{S}(2)$ the space of all $2\times 2$ symmetric positive definite matrices whose determinant is 1, then
@@
where $Q=(Q_{ij})\colon\Omega\rightarrow\mathbf{S}(2)$. Right multiplying both sides by $D_{f}(z)^{-1}$, we obtain a linear system which is the alternative formulation of the Beltami equation:
@@
and from this we have $\mu=\frac{q_{11}-q_{22}+2iq_{12}}{q_{11}+q_{22}+2\text{sgn}(\det{D_{f}})}$. With the constraint that $Q(z)\in\mathbf{S}(2)$, we observe that $|\mu(z)|<1$ if and only if $\det D_{f}(z)>0$. By inverse function theorem, researchers would through restricting $\|\mu\|_{\infty}<1$, achieve local bijectivity and prevent folding in the triangular mesh lui2014teichmuller; choi2015flash.
Suppose $\mu(z)$ is a measurable complex-valued function defined in a domain $U\subseteq\mathbb{C}$ for which $\|\mu\|_{\infty}<1$, we have the following existence theorem.
@@
(Measurable Riemann Mapping Theorem) For any function $\mu:U\to\mathbb{C}$ on with bounded essential supremum norm $\|\mu\|_{\infty}<1$, there is a quasiconformal map $\phi$ on $\overline{U}$ satisfying the Beltrami equation $\phi_{\overline{z}}=\mu\phi_{z}$ for almost all $z\in U$. Moreover, $\phi$ is unique up to post-composition with conformal isomorphisms and $\phi$ depends holomorphically on $\mu$.
When comes to discretization, it is a natural manner to ask whether there is a numerical method to compute the quasiconformal mapping given the BCs. The answer is yes, thanks to the Least Squares Quasiconformal Energy.
4.2 Least Squares Quasiconformal Energy
Apart from Eq.5, the Beltrami equation can be rewritten as
@@ 4.2 Least Squares Quasiconformal Energy
Here we assume $|\mu|<1$ and for convenience, we denote the square matrix on the right-hand side by $\mathbf{A}$. Inspired by the least squares conformal energy levy2002least, Qiu qiu2019computing proposed the concept least squares quasiconformal (LSQC) energy of a map.
@@
Let $\mu=\rho+ir$ be a complex-valued function defined on the domain $\Omega$. The LSQC energy of the map $z=(x,y)\mapsto(u,v)$ against the BC $\mu$ is defined to be
$E_{LSQC}(u,v,\mu)=\frac{1}{2}\int_{\Omega}\|P\nabla u+JP\nabla v\|^{2}\,dx\,dy,$ (7)
@@
$P=\frac{1}{\sqrt{1-|\mu|^{2}}}\begin{bmatrix}1-\rho&-r\\ -r&1+\rho\end{bmatrix},J=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$,
so that $P^{T}P=\mathbf{A}$ as in Eq.6 and $P^{T}JP=J$.
The LSQC energy admits the decomposition
$E_{LSQC}(u,v,\mu)=E_{\mathbf{A}}(u;\Omega)+E_{\mathbf{A}}(u;\Omega)-\mathcal{A}(u,v)$ (8)
where
$E_{\mathbf{A}}(u;\Omega)=\frac{1}{2}\int_{\Omega}\nabla u^{T}\mathbf{A}\nabla udxdy=\frac{1}{2}\int_{\Omega}\|P\nabla u\|^{2}dxdy$
is the Dirichlet-type energies of second-order equation $\nabla\cdot(\mathbf{A}\nabla u(z))=0$ derived by the decoupling method lui2014teichmuller and
$\mathcal{A}(u,v)=\int_{\Omega}(u_{y}v_{x}-u_{x}v_{y})dxdy.$ (9)
is the area of the image.
Given a piecewise linear map $f(x,y)=u+\sqrt{-1}v$ and consider a triangle $T=(x_{j},y_{j})_{j=1,2,3}$ of $\mathbb{R}^{2}$, we have
@@
where $d_{T}=(x_{1}y_{2}-y_{1}x_{2})+(x_{2}y_{3}-y_{2}x_{3})+(x_{3}y_{1}-y_{3}x_{1})$ is twice the area of the triangle. The Beltrami equation (Eq.3) can be rewritten as $(1-\mu)f_{x}+i(1+\mu)f_{y}=0$ and collected as a complex-valued linear system.
$0=\frac{i}{d_{T}}(W_{1}\quad W_{2}\quad W_{3})(U_{1}\quad U_{2}\quad U_{3})^{\top}$
@@
$W_{1}$ $=(1+\mu)(x_{3}-x_{2})+\sqrt{-1}(1-\mu)(y_{3}-y_{2})$
$W_{2}$ $=(1+\mu)(x_{1}-x_{3})+\sqrt{-1}(1-\mu)(y_{1}-y_{3})$
$W_{3}$ $=(1+\mu)(x_{2}-x_{1})+\sqrt{-1}(1-\mu)(y_{2}-y_{1}).$
@@
$E_{LSQC}(\textbf{U}=(U_{1},\ldots,U_{|\mathcal{V}|}))$ $=\sum\limits_{T\in\mathcal{T}}d_{T}\Big|\left((1-\mu)f_{x}+i(1+\mu)f_{y}\right)\Bigg|_{T}\Big|^{2}$
$=\sum_{T\in\mathcal{T}}\frac{1}{d_{T}}\Big|(W_{1,T}\qquad W_{2,T}\qquad W_{3,T})(U_{1,T}\qquad U_{2,T}\qquad U_{3,T})^{\top}\Big|^{2}$
and we may write as $E_{LSQC}(\textbf{U})=\|\mathcal{M}\textbf{U}\|^{2}$ with $\mathcal{M}=(m_{ij})\in\mathbb{C}^{|\mathcal{F}|\times|\mathcal{V}|}$ defined as
@@
To obtain a nontrivial solution, some of the $U_{i}$’s have to be pinned.
We decomposed $\mathbf{U}=(\mathbf{U}_{f}^{\top},\mathbf{U}_{p}^{\top})^{\top}$, where $\mathbf{U}_{f}$ are free points, i.e. variables of the optimization problem, and $\mathbf{U}_{f}$ are the points pinned. Similarly, we can decompose $\mathcal{M}$ in blocked matrices as
$\mathcal{M}=\begin{pmatrix}\mathcal{M}_{f},\mathcal{M}_{p}\end{pmatrix}$
with $\mathcal{M}_{f}\in\mathbb{C}^{|\mathcal{F}|\times(|\mathcal{V}|-p)}$ and $\mathcal{M}_{p}\in\mathbb{C}^{|\mathcal{F}|\times p}$. Let 1 and 2 be real and imaginary parts of a complex number, and rewrite the $|\mathcal{F}|$ complex linear equations to $2|\mathcal{F}|$ real equations, we have
$E_{LSQC}(\mathbf{U})=\|\mathcal{A}\mathbf{u}-\mathbf{b}\|^{2}$
@@
$\mathcal{A}=\begin{pmatrix}\mathcal{M}_{f}^{1}&-\mathcal{M}_{f}^{2}\\ \mathcal{M}_{f}^{2}&\mathcal{M}_{f}^{1}\end{pmatrix},\mathbf{b}=-\begin{pmatrix}\mathcal{M}_{p}^{1}&-\mathcal{M}_{p}^{2}\\ \mathcal{M}_{p}^{2}&\mathcal{M}_{p}^{1}\end{pmatrix}\begin{pmatrix}\mathbf{U}_{p}^{1}\\ \mathbf{U}_{p}^{2}\end{pmatrix}$.
Finally, the nontrivial solution is $\mathbf{U}=\begin{pmatrix}(\mathcal{A}^{\top}\mathcal{A})^{-1}\mathcal{A}^{\top}\mathbf{b}\\ \mathbf{U}_{p}\end{pmatrix}$.
Although LSQC has been used in numerical computations, its theoretical properties and suitability for neural simulation have not been systematically studied. We address this gap in the following section.
5 Properties of Least Squares Quasiconformal Energy
@@
In discretization, the numerical solution $U$ is unique when at least two points’ destinations are prescribed qiu2019computing. Here is the statement with proof.
@@
Suppose $\|\mu\|_{\infty}<1$, and the triangulation mesh is connected without dangling triangles. As long as $p\geq 2$, then $\mathcal{A}$ has full rank.
@@
The argument is essentially same as that in levy2002least. The idea is that start with the pinned points, we can construct the triangular mesh $\mathcal{T}$ incrementally via two operations:
@@
Correspondingly, the matrix $\mathcal{M}$ is constructed incrementally as well. Let $\mathcal{M}^{(i)}\in\mathbb{C}^{|\mathcal{F}_{i}|\times|\mathcal{V}_{i}|}$ be the $\mathcal{M}$ based on the constructed triangulation after $i$-th step. Two key observations are that $|\mathcal{F}_{i}|\geq|\mathcal{V}_{i}|-2$ and the rank of $\mathcal{M}_{f}^{(i)}$ is always its column number. If the $(i+1)$-th step is a joining operation, then compared to $\mathcal{M}_{f}^{(i)}$, $\mathcal{M}_{f}^{(i+1)}$ just has one more row, and joining has no affect on the rank. If the $(i+1)$-th step is a gluing operation, then $\mathcal{M}_{f}^{(i+1)}$ has one more row and column, and the new column has only one nonzero element in the last row thus the rank still equals to the column number. The premise of the argument is that for each triangle $T$, correspond $W_{j,T}$ are nonzero. This is still true when $\|\mu\|_{\infty}<1$. Let $\mu=\rho+\sqrt{-1}\tau$ and $L=\begin{pmatrix}1+\rho&\tau\\ \tau&1-\rho\end{pmatrix}$, then $W_{j,T}=\begin{pmatrix}1&\sqrt{-1}\end{pmatrix}L\begin{pmatrix}x_{j+2}-x_{j+1}\\ y_{j+2}-y_{j+1}\end{pmatrix}$. Since $\|\mu\|_{\infty}<1$ and $\mathcal{T}$ is a non-degenerate triangulation, then $L$ is non-singular and $W_{j,T}$ are all nonzero. Last, since $\mathcal{M}_{f}$ is a full-rank complex matrix, $\mathcal{A}$ also has full rank.
@@
Corollary 1
Suppose $\|\mu\|_{\infty}<1$ and face-wise constant, and $\mathbf{U}$ is the face-wise linear solution by solving the LSQC system with two points pinned. If $|\mathcal{F}|=|\mathcal{V}|-2$, then the BC for each face, induced by $\mathbf{U}$, is exactly the provided $\mu$.
Proof
Since the number of pinned points is 2 and no triangle face is predetermined, $\mathcal{M}_{f}$ has exactly $|\mathcal{F}|=|\mathcal{V}|-2$, equal to the number of columns. By above, we know $\mathcal{M}_{f}$ has full rank and thus $\mathcal{A}$ is an invertible square matrix. Therefore, $\mathbf{U}=\mathcal{A}^{-1}\mathbf{b}$ and $E_{\text{LSQC}}(\mathbf{U})=0$. Consequently, let $\mathbf{U}=u+\sqrt{-1}v.$, then $P\nabla u+JP\nabla v=0$ on each face.
However, it is more common that the LSQC energy functional of a given BC does not have a zero-value minimizer, because there is inherently no face-wise linear homeomorphism for the given face-wise constant BC or $|\mathcal{F}|\neq|\mathcal{V}|-2$. Despite this, once the provided face-wise constant BC is induced by a face-wise linear homeomorphism, the LSQC algorithm can reconstruct the homeomorphism given the induced BC and two points of the homeomorphism.
@@
With two points pinned, the LSQC system has only one minimizer satisfying the pin-point condition. Since $\mu$ is induced by $F$, $P\nabla u+JP\nabla v=0$ on each face and $E_{\text{LSQC}}(F)=0$. Meanwhile, $F$ satisfies the pin-point condition, and thus it is the solution to the LSQC system.
@@
In all, LSQC also has its own properties related to solution’s uniqueness and existence given points pinned, as well as invariant by a family of transformations, which is very analogous to Thm.4.1. Moreover, the solution of the minimization problem is independent of the resolution of the mesh.
@@
where $\mathcal{N}_{f}\in\mathbb{C}^{(|\mathcal{F}|-1)\times(|\mathcal{V}|-p)}$, $\mathcal{N}_{p}\in\mathbb{C}^{(|\mathcal{F}|-1)\times p}$, $\mathcal{F}\in\mathbb{C}^{1\times(|\mathcal{V}|-p)}$ with three nonzero elements $f_{j}=W_{j,T}/\sqrt{d_{T}}$, $\mathcal{L}$ and $\mathcal{P}$ both have three rows with nonzero elements defined as:

### v2

@@ 2.1 Quasiconformal Mappings
Quasiconformal mappings generalize conformal mappings by extending orientation-preserving homeomorphisms to those with bounded conformality distortions. Mathematically, suppose $\mathcal{U}$ is a domain in $\mathbb{C}$, a mapping $f\colon\mathcal{U}\rightarrow\mathbb{C}$ is a quasi-conformal map if it satisfies the Beltrami equation
@@
for some measurable complex-valued function $\mu$ with $\|\mu\|_{\infty}<1$, where $\frac{\partial f}{\partial z}=\frac{1}{2}(f_{x}-\sqrt{-1}f_{y})$ and $\frac{\partial f}{\partial\bar{z}}=\frac{1}{2}(f_{x}+\sqrt{-1}f_{y})$. $\mu$ is called the complex dilation or Beltrami coefficient (BC) of $f$, which measures the local deviation from a conformal map.
Let $f\colon(x,y)\rightarrow(u,v)$ and $\mu=\rho+\sqrt{-1}\tau$, and denote by $\mathbf{S}(2)$ the space of all $2\times 2$ symmetric positive definite matrices whose determinant is 1, then
@@
where $Q=(Q_{ij})\colon\Omega\rightarrow\mathbf{S}(2)$. Right multiplying both sides by $D_{f}(z)^{-1}$, we obtain a linear system which is the alternative formulation of the Beltrami equation:
@@
and from this we have $\mu=\frac{q_{11}-q_{22}+2\sqrt{-1}q_{12}}{q_{11}+q_{22}+2\text{sgn}(\det{D_{f}})}$. With the constraint that $Q(z)\in\mathbf{S}(2)$, we observe that $|\mu(z)|<1$ if and only if $\det D_{f}(z)>0$. By inverse function theorem, researchers can achieve local bijectivity and prevent folding by restricting $\|\mu\|_{\infty}<1$ in the triangular mesh Lui et al. (2014); Choi et al. (2015).
Suppose $\mu(z)$ is a measurable complex-valued function defined in a domain $\mathcal{U}\subseteq\mathbb{C}$ for which $\|\mu\|_{\infty}<1$, we have the following existence theorem.
@@
(Measurable Riemann Mapping Theorem) For any function $\mu:\mathcal{U}\to\mathbb{C}$ on with bounded essential supremum norm $\|\mu\|_{\infty}<1$, there is a quasiconformal map $\phi$ on $\overline{\mathcal{U}}$ satisfying the Beltrami equation $\phi_{\bar{z}}=\mu\phi_{z}$ for almost all $z\in\mathcal{U}$. Moreover, $\phi$ is unique up to post-composition with conformal isomorphisms and $\phi$ depends holomorphically on $\mu$.
When comes to discretization, it is a natural manner to ask whether there is a numerical method to compute the quasiconformal mapping given the BC. The answer is yes, thanks to the Least-squares Quasiconformal Energy.
Apart from Equation 7, the Beltrami equation can be rewritten as
@@
Here we assume $|\mu|<1$ and for convenience, we denote the square matrix on the right-hand side by $\mathbf{A}$. Inspired by the least-squares conformal energy Lévy et al. (2002), Qiu Qiu et al. (2019) proposed the concept least-squares quasiconformal (LSQC) energy of a map.
@@
Let $\mu=\rho+\sqrt{-1}r$ be a complex-valued function defined on the domain $\mathcal{U}$. The LSQC energy of the map $z=(x,y)\mapsto(u,v)$ against the BC $\mu$ is defined to be
$E_{LSQC}(u,v,\mu)=\frac{1}{2}\int_{\mathcal{U}}\|P\nabla u+JP\nabla v\|^{2}\,dx\,dy,$ (9)
@@
$P=\frac{1}{\sqrt{1-|\mu|^{2}}}\begin{bmatrix}1-\rho&-\tau\\ -\tau&1+\rho\end{bmatrix},J=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$,
so that $P^{T}P=\mathbf{A}$ as in Equation 8 and $P^{T}JP=J$.
The weight term $\frac{1}{\sqrt{1-|\mu|^{2}}}$ introduces numerical singularities as $|\mu|\to 1$ and renders the dependence of the energy on $\mu$ highly non-linear. Therefore, we adopt a simplified, unweighted formulation for the remainder of this work. We define the unweighted LSQC energy as the $\ell^{2}$-norm of the Beltrami equation mismatch.
Definition 2
Let $\mu=\rho+\sqrt{-1}\tau$ be a complex-valued function defined on the domain $\mathcal{U}$. The LSQC energy of the map $z=(x,y)\mapsto(u,v)$ against the BC $\mu$ is defined to be
$\tilde{E}_{LSQC}(f,\mu)=\frac{1}{2}\int_{\mathcal{U}}|f_{\bar{z}}-\mu f_{z}|^{2}dA.$
Equivalently,
$\tilde{E}_{LSQC}(u,v,\mu)=\frac{1}{2}\int_{\mathcal{U}}\|\tilde{P}\nabla u+J\tilde{P}\nabla v\|^{2}\,dx\,dy,$ (10)
where $\tilde{P}=\begin{bmatrix}1-\rho&-\tau\\ -\tau&1+\rho\end{bmatrix}$. Unless otherwise specified, all references to $E_{LSQC}$ in the subsequent sections refer to this unweighted formulation.
2.2 Discrete Setup and Notation for LSQC
Let $\mathcal{T}=(\mathcal{V},\mathcal{F})$ be a triangulation of a bounded planar domain $\mathcal{U}$, with vertex set $\mathcal{V}=\{1,\dots,N\}$, face (triangle) set $\mathcal{F}$. For each face $T\in\mathcal{F}$, we have vertices $p_{i}=(x_{i},y_{i})$, $i=1,2,3$, $d_{T}$ (twice the area), perimeter $h_{T}$ and the radius of the inscribed circle $r_{T}$. For the finite element space, we denote the Real P1 FE space:
$V_{h}:=\{\phi_{h}:\text{continuous, piecewise linear on }\mathcal{T}\}.$
with Nodal basis $\{\varphi_{i}\}_{i=1}^{N}\subset V_{h}$ with $\varphi_{i}(x_{j})=\delta_{ij}$. The Complex-valued P1 FE map is $f_{h}=u_{h}+\sqrt{-1}v_{h},\,u_{h},v_{h}\in V_{h}.$ and its full coefficient vector
$X(f_{h}):=(u_{1},\dots,u_{N},\,v_{1},\dots,v_{N})^{\top}\in\mathbb{R}^{2N}.$
As shown in the next section, to obtain a nontrivial LSQC minimizer, we need to pin at least two vertices in $\mathcal{T}$. We denote the pinned vertex index set by $P\subset\mathcal{V},\,|P|\geq 2.$ and prescribed images of pinned vertices, $q_{p}=(q_{p,x},q_{p,y})\in\mathbb{R}^{2},\,p\in P$. Thus, we need to introduce the zero-at-pins subspace in function and coefficient spaces,
$V_{h,0}:=\{f_{h}=u_{h}+iv_{h}:u_{h},v_{h}\in V_{h},\ f_{h}(p)=0\ \forall p\in P\},$
$\mathcal{X}_{0}:=\{X\in\mathbb{R}^{2N}:X_{p}=X_{N+p}=0\ \forall p\in P\}.$
and fixed pin map $f_{h,\mathrm{pin}}$ such that
$f_{h,\mathrm{pin}}(p)=q_{p}\quad(p\in P),\qquad f_{h,\mathrm{pin}}(v)=0\quad(v\notin P)$
with its full coefficient vector $\delta(P,q)\in\mathbb{R}^{2N}$:
$\delta_{j}:=\begin{cases}q_{p,x},&j=p,\ p\in P,\\[2.0pt] q_{p,y},&j=N+p,\ p\in P,\\[2.0pt] 0,&\text{otherwise}.\end{cases}$
Any total map satisfying pin constraints can be written as $f_{h}^{\mathrm{tot}}=f_{h}+f_{h,\mathrm{pin}},\,f_{h}\in V_{h,0},$ with $X(f_{h}^{\mathrm{tot}})=X(f_{h})+\delta(P,q).$ Let $\widehat{X}\in\mathbb{R}^{2N-2|P|}$ denote the coefficient vector at free vertices only, obtained from $X\in\mathcal{X}_{0}$ by deleting entries indexed by $p$ and $N+p$, $p\in P$, and correspondingly there is a fixed injection $\mathrm{lift}:\mathbb{R}^{2N-2|P|}\to\mathbb{R}^{2N}$ such that $X=\mathrm{lift}(\widehat{X})\in\mathcal{X}_{0},$ i.e. $\mathrm{lift}$ re-inserts zeros in the pinned coordinates.
In discretization, the BC means facewise constant Beltrami field
$\mu=(\mu_{T})_{T\in\mathcal{F}},\quad\mu_{T}=\rho_{T}+\sqrt{-1}\tau_{T},\quad\|\mu\|_{\infty}:=\max_{T\in\mathcal{F}}|\mu_{T}|.$
and denote $S:=\{\mu:\|\mu\|_{\infty}<1\}$. Given a piecewise linear map $f=u+\sqrt{-1}v$ and consider a triangle $T=(x_{j},y_{j})_{j=1,2,3}$ of $\mathbb{R}^{2}$, we have
@@ 2.2 Discrete Setup and Notation for LSQC
where $d_{T}=(x_{1}y_{2}-y_{1}x_{2})+(x_{2}y_{3}-y_{2}x_{3})+(x_{3}y_{1}-y_{3}x_{1})$ is twice the area of the triangle. The Beltrami equation (Equation 5) can be rewritten as $(1-\mu)f_{x}+\sqrt{-1}(1+\mu)f_{y}=0$ and collected as a complex-valued linear system.
$0=\frac{\sqrt{-1}}{d_{T}}(W_{1}\quad W_{2}\quad W_{3})(U_{1}\quad U_{2}\quad U_{3})^{\top}$
@@
$W_{1}$ $=(1+\mu)(x_{3}-x_{2})+\sqrt{-1}(1-\mu)(y_{3}-y_{2})$ (11)
$W_{2}$ $=(1+\mu)(x_{1}-x_{3})+\sqrt{-1}(1-\mu)(y_{1}-y_{3})$ (12)
$W_{3}$ $=(1+\mu)(x_{2}-x_{1})+\sqrt{-1}(1-\mu)(y_{2}-y_{1}).$ (13)
@@
$4\,E_{LSQC}(\textbf{U}=(U_{1},\ldots,U_{|\mathcal{V}|}))$ $=\sum\limits_{T\in\mathcal{T}}d_{T}\Big|\left((1-\mu)f_{x}+\sqrt{-1}(1+\mu)f_{y}\right)\Bigg|_{T}\Big|^{2}$
$=\sum_{T\in\mathcal{T}}\frac{1}{d_{T}}\Big|(W_{1,T}\quad W_{2,T}\quad W_{3,T})(U_{1,T}\quad U_{2,T}\quad U_{3,T})^{\top}\Big|^{2}$
and we may write as $4\,E_{LSQC}(\textbf{U})=\|\mathcal{M}\textbf{U}\|^{2}$ with $\mathcal{M}=(m_{ij})\in\mathbb{C}^{|\mathcal{F}|\times N}$ defined as
@@
Using the complex Beltrami equation on each face, we assemble $A_{\mathrm{full}}(\mu)\in\mathbb{R}^{2|\mathcal{F}|\times 2N}$ obtained by splitting real/imaginary parts of the complex system $\mathcal{M}\textbf{U}=0.$ Each row corresponds to a real equation; each row has only six nonzero vertex blocks. With pinned vertex index set $P$ given, we can obtain a ”reduced” LSQC matrix $A(\mu)$ by restricting the columns of $A_{\mathrm{full}}(\mu)$ to free vertices:
$A(\mu):=A_{\mathrm{full}}(\mu)\,\Pi\quad\in\mathbb{R}^{2|\mathcal{F}|\times(2N-2|P|)},$
where $\Pi:\mathbb{R}^{2N-2|P|}\to\mathbb{R}^{2N}$ is the matrix of the injection $\mathrm{lift}$. Thus,
$A_{\mathrm{full}}(\mu)\,\mathrm{lift}(\widehat{X})=A(\mu)\,\widehat{X}$.
Concretely, for a given $\mu$ and pinned point conditions $P,q$, the LSQC energy on total maps can be written in reduced variables as
$E_{LSQC}(\widehat{X};\mu):=\frac{1}{2}\big\|A(\mu)\widehat{X}+A_{\mathrm{full}}(\mu)\delta(P,q)\big\|_{2}^{2},$
and the LSQC minimizer $\widehat{X}(\mu)=\arg\min_{\widehat{X}}J(\widehat{X};\mu),$ is the unique solution of the normal equations
$A(\mu)^{T}A(\mu)\,\widehat{X}(\mu)=-A(\mu)^{T}A_{\mathrm{full}}(\mu)\,\delta(P,q).$
(non-singularity of $A(\mu)$ is shown in the next section). We define
$N(\mu):=A(\mu)^{T}A(\mu)\in\mathbb{R}^{(2N-2|P|)\times(2N-2|P|)}$
Concretely, We decomposed $\mathbf{U}=(\mathbf{U}_{f}^{\top},\mathbf{U}_{p}^{\top})^{\top}$, where $\mathbf{U}_{f}$ are free points, i.e. variables of the optimization problem, and $\mathbf{U}_{f}$ are the points pinned. Similarly, we can decompose $\mathcal{M}$ in blocked matrices as $\mathcal{M}=\begin{pmatrix}\mathcal{M}_{f},\mathcal{M}_{p}\end{pmatrix}$ with $\mathcal{M}_{f}\in\mathbb{C}^{|\mathcal{F}|\times(N-|P|)}$ and $\mathcal{M}_{p}\in\mathbb{C}^{|\mathcal{F}|\times|P|}$. Let 1 and 2 be real and imaginary parts of a complex number, and rewrite the $|\mathcal{F}|$ complex linear equations to $2\,|\mathcal{F}|$ real equations, we have
$E_{LSQC}(\mathbf{U};\mu)=\|A(\mu)\mathbf{u}-\mathbf{b}\|^{2}$
@@
$A(\mu)=\begin{pmatrix}\mathcal{M}_{f}^{1}&-\mathcal{M}_{f}^{2}\\ \mathcal{M}_{f}^{2}&\mathcal{M}_{f}^{1}\end{pmatrix},\mathbf{b}=-\begin{pmatrix}\mathcal{M}_{p}^{1}&-\mathcal{M}_{p}^{2}\\ \mathcal{M}_{p}^{2}&\mathcal{M}_{p}^{1}\end{pmatrix}\begin{pmatrix}\mathbf{U}_{p}^{1}\\ \mathbf{U}_{p}^{2}\end{pmatrix}$.
Although LSQC has been used in numerical computations, its theoretical properties, suitability for neural simulation and stability estimate have not been systematically studied. We address this gap in the following section.
3 Properties of Least-squares Quasiconformal Energy
@@
In discretization, the numerical solution $U$ is unique when at least two points’ destinations are prescribed. Here is the statement with proof.
@@
Suppose $\|\mu\|_{\infty}<1$, and the triangulation mesh is connected without dangling triangles. As long as $p\geq 2$, then $\mathcal{A}=A(\mu)$ has full rank.
@@
The idea is that start with the pinned points, we can construct the triangular mesh $\mathcal{T}$ incrementally via two operations:
@@
Correspondingly, the matrix $\mathcal{M}$ is constructed incrementally as well. Let $\mathcal{M}^{(i)}\in\mathbb{C}^{|\mathcal{F}_{i}|\times|\mathcal{V}_{i}|}$ be the $\mathcal{M}$ based on the constructed triangulation after $i$-th step. Two key observations are that $|\mathcal{F}_{i}|\geq|\mathcal{V}_{i}|-2$ and the rank of $\mathcal{M}_{f}^{(i)}$ is always its column number. If the $(i+1)$-th step is a joining operation, then compared to $\mathcal{M}_{f}^{(i)}$, $\mathcal{M}_{f}^{(i+1)}$ just has one more row, and joining has no affect on the rank. If the $(i+1)$-th step is a gluing operation, then $\mathcal{M}_{f}^{(i+1)}$ has one more row and column, and the new column has only one nonzero element in the last row thus the rank still equals to the column number. The premise of the argument is that for each triangle $T$, correspond $W_{j,T}$ are nonzero. This is still true when $\|\mu\|_{\infty}<1$. Let $\mu=\rho+\sqrt{-1}\tau$ and $G(\mu_{T})=\begin{pmatrix}1+\rho_{T}&\tau_{T}\\ \tau_{T}&1-\rho_{T}\end{pmatrix}$, then $W_{j,T}=\begin{pmatrix}1&\sqrt{-1}\end{pmatrix}G(\mu_{T})\begin{pmatrix}x_{j+2}-x_{j+1}\\ y_{j+2}-y_{j+1}\end{pmatrix}$. Since $\|\mu\|_{\infty}<1$ and $\mathcal{T}$ is a non-degenerate triangulation, then $G(\mu_{T})$ is non-singular and $W_{j,T}$ are all nonzero. Last, since $\mathcal{M}_{f}$ is a full-rank complex matrix, $\mathcal{A}$ also has full rank.
@@
Besides, suppose the provided face-wise constant BC is induced by a face-wise linear homeomorphism, the LSQC algorithm can reconstruct the homeomorphism given the induced BC and two points of the homeomorphism.
@@
With two points pinned, the LSQC system has only one minimizer satisfying the pin-point condition. Since $\mu$ is induced by $F$, $(1-\mu)F_{x}+\sqrt{-1}(1+\mu)F_{y}=0$ on each face and $E_{\text{LSQC}}(F)=0$. Meanwhile, $F$ satisfies the pin-point condition, and thus it is the solution to the LSQC system.
@@
In all, LSQC also has its own properties related to solution’s uniqueness and existence given points pinned, as well as invariant by a family of transformations, which is very analogous to Theorem 2.1. Moreover, the solution of the minimization problem is independent of the resolution of the mesh.
@@
$\mathcal{M}_{p}=\left(\begin{array}[]{ccc}&&\\ &\mathcal{N}_{p}&\\ &&\\ \hline\cr&\mathbf{0}&\end{array}\right),\quad\mathcal{M}_{p}^{+}=\left(\begin{array}[]{ccc}&&\\ &\mathcal{N}_{p}&\\ &&\\ \hline\cr&\mathbf{0}&\\ &\mathbf{0}&\\ &\mathbf{0}&\\ \end{array}\right).$
where $\mathcal{N}_{f}\in\mathbb{C}^{(|\mathcal{F}|-1)\times(|\mathcal{V}|-|P|)}$, $\mathcal{N}_{p}\in\mathbb{C}^{(|\mathcal{F}|-1)\times|P|}$, $\mathcal{F}\in\mathbb{C}^{1\times(|\mathcal{V}|-|P|)}$ with three nonzero elements $f_{j}=W_{j,T}/\sqrt{d_{T}}$, $\mathcal{L}$ and $\mathcal{P}$ both have three rows with nonzero elements defined as:
@@
$\mathcal{M}_{p}=\left(\begin{array}[]{ccc}&&\\ &\mathcal{N}_{p}&\\ &&\\ \hline\cr\mathbf{0}&f_{1}&\mathbf{0}\end{array}\right),\quad\mathcal{M}_{p}^{+}=\left(\begin{array}[]{ccc}&&\\ &\mathcal{N}_{p}&\\ &&\\ \hline\cr&\mathbf{0}&\\ \mathbf{0}&l_{21}&\mathbf{0}\\ \mathbf{0}&l_{31}&\mathbf{0}\\ \end{array}\right).$